Students sold 525 raffle tickets to raise money only 1 ticket will be drawn and 1 prize given away sergio brought 15 tickets wha is the probabilty that sergio will not win
Answer by stanbon(75887)

Students sold 525 raffle tickets to raise money only 1 ticket will be drawn and 1 prize given away sergio brought 15 tickets what is the probabilty that sergio will not win.
---------------------
Binomial with n = 15, p = 1/525, x = 0
--------------------------
P(x=0) = 15C0 (1/525)^0 (524/525)^15 = 0.9718..
Cheers,
Stan H.